OpenStudy (anonymous):

Show that Ek = 1/2mv2 is to respect base units

OpenStudy (anonymous):

unit of Energy=kgm^2/s^2 unit of mass=kg unit of velocity=m/s then unit of energy from 1/2 mv^2 will be=kgm^2/s^2

OpenStudy (anonymous):

thanks